OpenStudy (samsan9):

find the first 5 terms of the sequence given by a1=2, an=3an-1-1

OpenStudy (anonymous):

a1=2 a2=3a1 - 1=3*2 -1 =5 a3=3a2 -1 =3*5 -1 = 14 and so on..

OpenStudy (samsan9):